Problem 30-3 Z-Score

Fair-to-Midland Manufacturing, Inc. (FMM), has applied for a loan at True Credit Bank. Jon Fulkerson, the credit analyst at the bank, has gathered the following information from the company’s financial statements:

  



  Total assets$115,000 
  EBIT7,900 
  Net working capital4,100 
  Book value of equity23,000 
  Accumulated retained earnings19,600 
  Sales112,000 

   

The stock price of FMM is $31 per share and there are 7,800 shares outstanding. What is the Z-score for this company? (Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.)
Answer #1

Z-score = 1.2A + 1.4B + 3.3C + 0.6D + 1.0E


Where,


A = Working capital / Total Assets


B = Retained earnings / Total Assets


C = Earnings Before Interest and Tax / Total Assets


D = Market value of equity / Total Liabilities


E = Sales / Total Assets


Z-score = 1.2 \times ($4,100 / $115,000) + 1.4 \times ($19,600 / $115,000) + 3.3 \times ($7,900 / $115,000) + 0.6 \times ($241,800 / $92,000) + 1.0 \times ($112,000 / $115,000).


Z-score = 0.04278 + 0.23861 + 0.2267 + 1.5770 + 0.97391


Z-score = 3.059


Total liabilities = Total assets - Book value of equity


Total liabilities = $115,000 - $23,000 = $92,000.


answered by: gavin